OpenStudy (ccastudentgrade10):

help!!! what is the solution of the following system? {-2x - y = 1 {-4x - 2y = -1 a. infinite solutions b. no solutions c. (3, 8) d. (-3, -8) i am pretty sure it isnt c or d

OpenStudy (ccastudentgrade10):

@Vocaloid

OpenStudy (texaschic101):

-2x - y = 1 ..... -2x - 1 = y -4x - 2y = -1 -4x - 2(-2x - 1) = -1 -4x + 4x + 2 = -1 2 = -1 (incorrect)....this has no solutions

OpenStudy (ccastudentgrade10):

thank you! @texaschic101

OpenStudy (texaschic101):

sure thing :)